OpenStudy (aaronandyson):

if sin(theta)sec(theta) = -1 and theta lies in second quadrant ,find sin theta and sec theta

jimthompson5910 (jim_thompson5910):

http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf look at the `Reciprocal Identities` section (page 2) and you'll see that \[\Large \sec(\theta) = \frac{1}{\cos(\theta)}\]

jimthompson5910 (jim_thompson5910):

also look in the `Tangent and Cotangent Identities` section just above the reciprocal section and you'll see \[\Large \tan(\theta) = \frac{\sin(\theta)}{\cos(\theta)}\]

jimthompson5910 (jim_thompson5910):

Doing all this will hopefully convince you that \[\Large \sin(\theta)\sec(\theta) = -1\] is the same as \[\Large \tan(\theta) = -1\]

OpenStudy (aaronandyson):

okay, what next?

OpenStudy (aaronandyson):

Then tan theta = -1/1

jimthompson5910 (jim_thompson5910):

Draw an xy axis 

jimthompson5910 (jim_thompson5910):

we're in quadrant 2 now